OpenStudy (anonymous):

Multiply. ^3sqrt7 * ^3sqrt49 is the answer ^3sqrt3?

OpenStudy (anonymous):

no the answer is 7

OpenStudy (anonymous):

you get \[\sqrt[3]{7\times 9}=\sqrt[3]{7^3}=7\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

you have an unlimited supply of these?

OpenStudy (anonymous):

seems like it